1200kg car is traveling at a velocity of 20m/s the car then does 3000W of power for 20 seconds what is the change in momentum?


1
Expert's answer
2021-05-06T07:41:41-0400

Work done by car

"W=3000\times20=60000J"


Work done will be equal to change my n kinetic energy


"\frac{1}{2}(1200) (v^2-20^2) =60000\\v=22.36m/s"


Change in momentum

"\Delta P=1200(22.36-20) =2832kgm/s^2"